OpenStudy (anonymous):

A bag has 4 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a red marble, do not replace it, and pick another red marble?

OpenStudy (tkhunny):

\(\dfrac{Looking\;For}{All\;Possibilities} = \dfrac{Red}{Everything}\cdot\dfrac{Red}{Everything\;less\;One\;Red}\)

OpenStudy (anonymous):

you add 4 3 and 3=10.. then right down 3/10. after that, subtract 10 and one= 9 easy. know it says to do not replace the marble... so you have 4 green, 2 red, amd 3 yellow. so now you have 2/9... multiply 3/10 and 2/9... then tada!!! your answer